Good Evening,
I'm having a hard time catching on to linear equations and graphing them. Can you please show me how to figure out the answer for this problem.
5X-Y=-8
Thank you in advance for your help.

Answer by mananth(16949)

5x-y=-8
y=5x+8
..
Find the co-ordinates of the intercepts
y=0, x=-8/5 (-8/5,0)
x=0,y=8 (0,8)
Plot these points on the graph and draw the line joining the two points.